OpenStudy (anonymous):

What is the value of the expression 1 – r – a when r = 20 and a = 6? A. –27 B. –25 C. –15 D. –13

OpenStudy (anonymous):

challenge :(((

OpenStudy (bishoppatton):

B. -25

OpenStudy (anonymous):

i thought it was -13

OpenStudy (bishoppatton):

ONE SEC

OpenStudy (bishoppatton):

No because. 1 - 20 = -19 -19 - 6 = -25

OpenStudy (anonymous):

oh

OpenStudy (bishoppatton):

Negative - Positive = Negative

OpenStudy (anonymous):

oh ok

OpenStudy (bishoppatton):

Positive - Negative = Positive 9 - -6 = 15

OpenStudy (bishoppatton):

So subtracting negative from positive is kinda like doing addition with that -6 and turning it into a positive 6 so all you do is add 9 + 6 = 15

OpenStudy (bishoppatton):

Am i making sense to you?

OpenStudy (anonymous):

so negative and negative is positive positive and positve is positve negative and positive is negative positive and negative is positve

OpenStudy (bishoppatton):

Yes!!! Nice Job.

OpenStudy (anonymous):

ok thx but i thought it was wrong cuz 3 positve and 1 negative is wierd

OpenStudy (bishoppatton):

Write me a testimonial if i helped you learn something here. I know it seems weird but that is what it is!!!

OpenStudy (bishoppatton):

It is a good thing that you are learning as you work on this website. Good Job again man :)

OpenStudy (bishoppatton):

ok... i will brb

OpenStudy (anonymous):

ok

OpenStudy (bishoppatton):

i am back